OpenStudy (anonymous):

Please help SALES: The rate of increase in sales S (in thousands of units) of a product is proportional to the current level of sales and inversely proportional to the square of the time t. This is described by the differential equation: dS / dt = kS / t^2 where t is the time in years. The saturation point for the market is 50,000 units. That is, the limit as t approaches infinity is 50. After 1 year, 10,000 units have been sold. Find S as a function of t.

OpenStudy (anonymous):

idk im sorry

OpenStudy (anonymous):

its all good, i have no idea whats going on

OpenStudy (anonymous):

You have: dS/dt = k*S/(t^2) As you say, this is a separable equation: dS/S = k/(t^2) dt Integrate both sides: ln(S) - ln(c) = -k/t where ln(c) is the constant of integration. ln(S/c) = -k/t S(t) = c*exp(-k/t) As t -> ∞, exp(-k/t) -> 1, so lim(t -> ∞) S(t) = c = 50,000 So: S(t) = 50,000*exp(-k/t) We are also told that at t = 1 yr, S(1 yr) = 10,000. We can use this to solve for k: 10,000 = 50,000*exp(-k/yr) 1/5 = exp(-k/yr) ln(1/5) = -k/yr -k = ln(1/5)*yr ~= -1.609*yr So: S(t) = 50,000*exp(ln(1/5)*yr/t) ~= 50,000*exp((-1.609*yr)/t)

OpenStudy (anonymous):

thx alot im gonna examine this

OpenStudy (anonymous):

ok